1) A calorimeter contains 28.0 mL of water at 11.5 ?C . When 2.20 g of X (a substance with a molar mass of 51.0 g/mol ) is added, it dissolves via the reaction

X(s)+H2O(l)?X(aq)

and the temperature of the solution increases to 30.0 ?C .

Calculate the enthalpy change, ?H, for this reaction per mole of X.

Assume that the specific heat of the resulting solution is equal to that of water [4.18 J/(g??C)], that density of water is 1.00 g/mL, and that no heat is lost to the calorimeter itself, nor to the surroundings.

2)

Consider the reaction

C12H22O11(s)+12O2(g)?12CO2(g)+11H2O(l)

in which 10.0 g of sucrose, C12H22O11, was burned in a bomb calorimeter with a heat capacity of 7.50 kJ/?C. The temperature increase inside the calorimeter was found to be 22.0 ?C. Calculate the change in internal energy, ?E, for this reaction per mole of sucrose.
Answer #1

1.

Heat change = mass * specifc heat * change in temperature

q = (28.0 + 2.20) * 4.18 * (30.0 - 11.5)

q = 2335.4 J

Moles of X = mass / molar mass = 2.20 / 51.0 = 0.0431 mol

deltaHsol = - q / n = - 2335.4 / 0.0431 = - 54139 J / mol = - 54.1 kJ/mol

2.

q = Ccal * chang in temperature

q = 7.50 * 22.0

q = 165. kJ

Moles of sucrose = mass / molar mass = 10.0 / 342 = 0.0293 mol

deltaErxn = - q / n = - 165 / 0.0293 = - 5643 kJ/mol